OpenStudy (anonymous):

find the derivative of 6 sqrtx + 3 ^3sqrtx ?

OpenStudy (anonymous):

try changing the square roots to exponents, then apply power rule

OpenStudy (baldymcgee6):

okay, so we need to do some 'get ready steps'

OpenStudy (anonymous):

for example, sqrt(x) = x^(1/2)

OpenStudy (baldymcgee6):

\[h(x) =6\sqrt{x} + 3\sqrt[3]{x} = 6(x)^{1/2} + 3(x)^{1/3}\]

OpenStudy (anonymous):

what rule is this? @baldymcgee6

OpenStudy (anonymous):

rule of exponents

OpenStudy (baldymcgee6):

we haven't applied any differentiating rules yet. this is just re-arranging to make it easier for us.

OpenStudy (anonymous):

ohh ok ok

OpenStudy (anonymous):

so after that we pick rule right.

OpenStudy (baldymcgee6):

Right, here we can just use the power rule again

OpenStudy (anonymous):

ok so it would be 6x^1/2?

OpenStudy (anonymous):

@baldymcgee6

OpenStudy (anonymous):

No it would be